OpenStudy (anonymous):

Find the vertex, the line of symmetry the maximum or minimum value of the quadratic function, and graph the function f(x)3x^2-12x+12

OpenStudy (anonymous):

3(x-2)(x-2)

OpenStudy (anonymous):

use x = -b/2a

OpenStudy (anonymous):

a = 3, b = -12

OpenStudy (anonymous):

equation for vertex(maximum/minimum). If u have a quadratic in the form of y=ax^2+bx+c then the vertex is when x=-b/2a

OpenStudy (anonymous):

(2,0)

OpenStudy (anonymous):

This would be your minimum value

OpenStudy (anonymous):

x = 2 is line of symmetry

OpenStudy (anonymous):

in this case a=3 b=-12 c=12 so using the equation: x=-(-12)/(2*3)=2 so solve for the y value

OpenStudy (anonymous):

I just needed the first part thanks